### v1

# Probing Pre-trained Language Models on Code Changes: Insights from ReDef, a High-Confidence Just-in-Time Defect Prediction Dataset

Abstract.
Just-in-Time software defect prediction (JIT-SDP) plays a critical role in prioritizing risky code changes during code review and continuous integration. However, existing datasets often suffer from noisy labels and low precision in identifying bug-inducing commits. To address this, we present ReDef (Revert-based Defect dataset), a high-confidence benchmark of function-level modifications curated from 22 large-scale C/C++ projects. Defective cases are anchored by revert commits, while clean cases are validated through post-hoc history checks. Ambiguous instances are conservatively filtered out via a GPT-assisted triage process involving multiple votes and audits. This pipeline yields 3,164 defective and 10,268 clean modifications, offering substantially more reliable labels than prior existing resources. Beyond dataset construction, we provide the first systematic evaluation of how pre-trained language models (PLMs) reason about code modifications—specifically, which input encodings most effectively expose change information, and whether models genuinely capture edit semantics. We fine-tune CodeBERT, CodeT5+, and UniXcoder under five encoding strategies, and further probe their sensitivity through counterfactual perturbations that swap added/deleted blocks, invert diff polarity, or inject spurious markers. Our results show that compact diff-style encodings consistently outperform whole-function formats across all PLMs, with statistical tests confirming large, model-independent effects. However, under counterfactual tests, performance degrades little or not at all—revealing that what appears to be robustness in fact reflects reliance on superficial cues rather than true semantic understanding. These findings indicate that, unlike in snapshot-based tasks, current PLMs remain limited in their ability to genuinely comprehend code modifications.
1. Introduction
Software defect prediction (SDP) aims to estimate whether a given piece of code is likely to contain defects so that limited engineering resources—testing, review time, and Continuous Integration (CI) capacity—can be prioritized where they matter most (Malhotra, 2015). Unlike defect detection or localization, which seeks to identify concrete bug instances, SDP is a risk estimation problem used to triage changes and guide quality assurance activities. Effective SDP reduces downstream failures and lowers maintenance costs by surfacing risky code early enough for developers to take action within their normal workflow—for example, during code review or before a merge in CI.
Classical SDP typically models static snapshots of code or aggregates hand-crafted process metrics—such as churn, complexity, ownership, and past defects—at the file, module, or function level (Wahono, 2015). While such predictors can correlate with risk, they fail to capture how the code is evolving at the moment of change. Just-in-Time (JIT) SDP addresses this limitation by reasoning over code modifications at commit time, aligning closely with practical review and CI processes that scrutinize the changed lines (Kamei et al., 2012; Kim et al., 2008). However, most existing JIT-SDP datasets rely on SZZ(Śliwerski–Zimmermann–Zeller)-style linkages from bug-fixing commits back to bug-introducing commits (Śliwerski et al., 2005). SZZ attributes the introduction of a defect to the last modifier of the lines changed by a later fix—an assumption that is often violated by routine edits such as renaming or refactoring. This approach introduces substantial noise: it can misidentify the true introduction points, propagate commit-level label errors, and mismatch the actual defective lines, ultimately obscuring evaluation quality (Herbold et al., 2022; Lyu et al., 2024). To mitigate the noise inherent in SZZ-based labeling, we turn to a more explicit and high-confidence signal: revert commits.
Reverts are a routine part of disciplined software development. Developers may revert commits for various reasons—such as regressions, incompatibilities, performance degradations, policy or formatting violations, or simply because another change has superseded the patch (Yan et al., 2019). A revert creates an explicit, time-stamped link to the earlier change. When the revert message cites a failure or fix, it provides strong evidence that the prior modification was defective. We therefore use revert commits as anchors to harvest high-confidence positive examples, and subsequently filter out cases unrelated to defective modifications.
With a more reliable labeling strategy in place, the next question is how modern models can be effectively leveraged. Existing JIT-SDP research has largely depended on compact engineered features with classical machine learning (ML) and, more recently, deep learning (DL) models. However, these approaches primarily capture surface-level correlations—such as churn, ownership, and developer activity—rather than the true semantics of code changes (Kamei et al., 2012, 2016; Hoang et al., 2019). In other words, they estimate probabilities from shallow cues instead of reasoning about what the code actually does.
By contrast, recent code-focused pre-trained language models (PLMs) and large language models (LLMs) have demonstrated strong performance across diverse SE tasks—such as code search, summarization, repair, and vulnerability detection—precisely because they are designed to capture the meaning of code itself, not merely its statistical correlates (Guo et al., 2022; Xia et al., 2023; Ni et al., 2023). This raises a natural question for JIT-SDP: if PLMs can effectively understand static code snapshots, can they also reason about code modifications, i.e., the relationship between the before and after states? Given that PLMs have consistently shown strong effectiveness for code understanding tasks, we therefore focus our study on PLMs as the primary models for systematic evaluation (Zhang et al., 2025b; Sun et al., 2023; Zhang et al., 2025a).
To answer this question, we first construct a high-confidence, change-centric benchmark. Next, in Stage 1, we compare alternative ways of presenting modification information to PLMs, asking which input encodings yield the most effective performance. For this, we fine-tune three widely used code PLMs—CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023b), and UniXcoder (Guo et al., 2022)—as backbones. However, strong results under these encodings alone do not guarantee genuine semantic reasoning, as models may still rely on superficial cues. Therefore, in Stage 2, we deliberately perturb the representation of code changes—e.g., by swapping added and deleted code blocks, inverting edit polarity, or injecting spurious markers—to test whether PLMs degrade when true semantic information is distorted. By probing how models behave under such distortions, our protocol not only measures raw accuracy but also uncovers a blind spot in current PLM research: apparent robustness may simply reflect reliance on shallow cues rather than genuine semantic reasoning about code changes.
Our key contributions are as follows:
High-Confidence JIT-SDP Dataset (ReDef): We curate ReDef (Revert-based Defect dataset), a corpus of single-function changes harvested via revert anchors and refined through GPT-assisted filtering. The dataset is further validated through manual sampling to ensure benchmark reliability, and its 10K+ (3,164 defective, 10,268 clean) scale makes it well-suited for fine-tuning PLMs.
Encoding Strategies for PLMs: We conduct the first systematic comparative study of five encoding strategies for representing code modifications as PLM inputs, providing clear empirical evidence of how input representations affect change-centric defect prediction.
Counterfactual Robustness Analysis: We design counterfactual perturbation tests to probe whether PLMs truly reason about code changes beyond surface cues, revealing limitations of current models and offering guidance for future research.
Together, these contributions provide both a reliable benchmark and empirical insights into how PLMs handle code changes, informing future defect prediction research and practice. The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 describes our dataset construction and PLM input strategies. Section 4 outlines the experimental setup, and Section 5 reports results. Section 6 discusses threats to validity, and Section 7 concludes.
2. Background and Related Work
This section reviews prior research on software defect prediction from three perspectives. Section 2.1 outlines classical Software Defect Prediction (SDP), which predicts defect-prone modules from static snapshots using either hand-crafted metrics or learned representations. Section 2.2 covers Just-in-Time software defect prediction (JIT-SDP), which shifts the unit of analysis to commits or patches and incorporates both engineered features and recent neural approaches. It also summarizes the limitations of existing JIT-SDP datasets, such as noise introduced by SZZ-based labeling. Finally, Section 2.3 discusses advances in applying Pre-trained Language Models (PLMs) across software engineering (SE) tasks, motivating their use for change-level defect prediction.
2.1. Software Defect Prediction (SDP)
Classical SDP focuses on artifact-level prediction from static snapshots, typically at the file, class, or module granularity. Early approaches rely on hand-crafted product metrics (e.g., size, complexity, coupling) and process/organization metrics (e.g., churn, revisions, developer ownership, past defects) (Malhotra, 2015; Bell et al., 2013; Huda et al., 2017). These features are commonly fed into traditional learners such as logistic regression, random forests, and Support Vector Machines (SVMs). For example, Elish et al. showed that SVMs perform competitively on NASA datasets compared to logistic regression (Elish and Elish, 2008), while ensemble methods based on random forests also address class imbalance (Balaram and Vasundra, 2022). Decision tree–based models have similarly achieved strong performance (Wang et al., 2012; Khoshgoftaar and Seliya, 2002). More recent work replaces hand-crafted metrics with learned representations, leveraging topic models (Nguyen et al., 2011), deep generative models (Wang et al., 2018), or transformer-based encoders (Omri and Sinz, 2020). Evaluation often considers effort-aware measures (e.g., recall at top-x% LOC) and cross-project transferability (Bell et al., 2013; Ostrand and Weyuker, 2007; Tosun et al., 2009; Turhan et al., 2009). Despite this progress, snapshot-based prediction cannot capture the semantics of specific changes, and file- or module-level labels blur defective regions—complicating causal interpretation and limiting actionable insights for change-level prediction.
2.2. Just-in-Time Software Defect Prediction (JIT-SDP) and Datasets
JIT-SDP shifts the focus from static artifacts to change-time risk estimation, scoring each commit or patch so that reviewers and CI gates can provide immediate feedback. The concept originates from Kim et al., who first classified individual changes as clean or buggy (Kim et al., 2008). Subsequent work engineered compact commit-level features—such as churn, diffusion, prior defect density, developer experience, recency, and message cues—and applied machine learning techniques (Kamei et al., 2012; Shihab et al., 2012; Kamei et al., 2016). Later studies incorporated commit messages and patch text, while neural models began encoding diffs directly. For example, Yang et al. proposed Deeper, a Deep Belief Network (DBN)-based model that outperformed logistic regression (Yang et al., 2015), and Hoang et al. introduced DeepJIT, which integrates commit messages and code metrics (Hoang et al., 2019). CC2Vec further embedded code changes hierarchically to capture structural and sequential information (Hoang et al., 2020).
More recently, researchers have applied PLMs to JIT-SDP. Guo et al. systematically compared six transformer-based backbones (RoBERTa, CodeBERT, BART, PLBART, GPT2, and CodeGPT) and showed that PLM-based variants consistently outperform traditional baselines such as DeepJIT and CC2Vec (Guo et al., 2023). Building on this, Abu Talib et al. demonstrated that parameter-efficient fine-tuning (PEFT) techniques such as LoRA can further boost JIT-SDP performance on ApacheJIT (Abu Talib et al., 2024), while Lin et al. introduced CCT5—a CodeT5-initialized model pre-trained on CodeChangeNet—that achieves state-of-the-art results on code-change tasks including JIT-SDP (Lin et al., 2023). Jiang et al. extended CodeBERT with a Replaced Message Identification objective, resulting in BiCC-BERT and JIT-BiCC, which improved F1-scores by over 10% (Jiang et al., 2025). Together, these studies highlight the promise of PLMs for JIT-SDP, yet it remains unclear whether current gains reflect genuine modeling of code-change semantics or incidental signals from existing datasets—underscoring the need for both more reliable benchmarks and systematic evaluation of model behavior.
Despite these advances, a critical limitation of existing JIT-SDP datasets lies in label reliability. Afric et al. reported that even with RoBERTa-based issue classification, file-level datasets contained on average 14.3% mislabeled instances (Afric et al., 2023). Herbold et al. showed that SZZ-based labeling is particularly error-prone, with only about half of the commits identified as bug-inducing being correct (Herbold et al., 2022). Several variants of SZZ attempt to mitigate this noise—for example, by excluding refactoring edits that standard SZZ mistakenly flags as defect-inducing or by tracing commit chains more accurately to recover missed bug-introducing links (Niu et al., 2025; Lyu et al., 2024). However, even with such refinements, existing datasets still fall short in reliability, underscoring the need for explicit validation of defect labels, which we address in RQ1.
2.3. Pre-trained Language Models (PLMs) in Other Software Engineering Tasks
Pre-trained language models (PLMs) trained on large code corpora—such as CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5—have been increasingly applied to diverse SE tasks, often achieving state-of-the-art results.
In code review, Li et al. proposed CodeReviewer, a Transformer-based encoder–decoder PLM tailored for diff review, improving both comment generation and code refinement quality (Li et al., 2022). Similarly, Kartal et al. demonstrated that transformer-based similarity measures outperform prior baselines in automated comment recommendation (Kartal et al., 2024).
In program repair, PLMs have achieved significant progress. For instance, Mashhadi and Hemmati applied CodeBERT to repair Java bugs, achieving 19–72% fix accuracy on ManySStuBs4J datasets (Mashhadi and Hemmati, 2021). Retrieval-augmented methods like RAP-Gen combine CodeT5 with external bug-fix retrieval, showing marked improvements over previous approaches—repairing 15 more bugs on the Defects4J benchmark (Wang et al., 2023a). Additionally, AlphaRepair uses CodeBERT in a zero-shot, infilling-style repair setup and achieves state-of-the-art results without task-specific retraining (Xia et al., 2023).
In vulnerability detection, PLMs have shown strong effectiveness across diverse settings. For example, one study fused multi-modal knowledge (structural, semantic, and syntactic features) with PLM encoders to achieve state-of-the-art function-level vulnerability detection (Ni et al., 2023). Atiiq et al. fine-tuned PLMs on the multi-language CVEFixes dataset (JavaScript, Java, Python, PHP, Go, C/C++), reporting substantial variation across languages and finding that UniXcoder outperformed other models (Atiiq et al., 2024). More recently, a comprehensive benchmark demonstrated that PLMs can surpass much larger code LLMs in accuracy and F1-scores, underscoring their superior inductive bias and efficiency for vulnerability detection tasks (Zhang et al., 2025b).
Taken together, PLMs have shown impressive capabilities across diverse SE tasks. Importantly, many of these tasks—such as program repair and vulnerability detection—require reasoning about the semantics of code edits or risky patterns, which parallels the demands of defect prediction. Their strong performance in such semantics-sensitive settings suggests that PLMs are particularly promising for change-aware defect prediction, motivating our systematic investigation in this paper.
3. Methodology
Our methodology comprises two main pillars. (A) ReDef Construction: We curate defective and clean code modifications from large-scale C/C++ projects, leveraging revert signals, post-hoc history screening, and expert audits to ensure high label reliability. (B) PLM Evaluation Protocol: We design a two-stage framework to assess how models interpret code modifications. Figure 1 summarizes the overall approach.
3.1. ReDef Construction
We curated ReDef, a high-confidence JIT-SDP corpus from 22 long-lived, high-activity C/C++ repositories. Projects were selected for sustained commit volume, explicit revert workflows, and domain diversity to reduce project-specific bias. The curation yielded 3,164 defective and 10,268 clean modifications. Figure 2 reports per-project counts.
Defective Modification. We extracted commits that touched exactly one file and whose modified lines were confined to a single function. This design ensures reliable causal attribution to a specific function—a choice also adopted in recent datasets such as PrimeVul (Ding et al., 2024). From this single-function pool, we identified defective modifications by collecting revert commits whose messages (i) contained the keyword “revert” and (ii) explicitly referenced the hash of an earlier commit. Since such reverts typically indicate that the referenced commit introduced a problem, the targeted commits are highly likely to represent defective modification. For each referenced hash, we retrieved its associated commit (the target of the revert) and collected four artifacts for analysis: the reverted commit message, the function_before and function_after from that commit, and the revert commit message itself. Together these formed an evidence bundle and were submitted to GPT-4o three independent times under a fixed rubric to remove ambiguous or non-defect reverts (e.g., formatting changes, policy reversions, or hash-only messages without rationale). The rubric distinguished each revert as: (A) Failed/incomplete fix (DefectiveDefective), (B) Bug introduction (CleanDefective), or (C) Other (non-defect). Categories (A) and (B) were merged into the defective class, while (C) served as a filter to discard non-defect cases. To ensure high precision, a candidate was retained only if all three GPT-4o votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B). Any sample with one or more (C) votes (e.g., A/A/C) was conservatively discarded.
Clean Modification. To assemble the clean stratum, we first identified all defective modifications and then targeted clean modification commits from a similar time period in the same repository, thereby mitigating both project-specific and period-specific bias. As with the defective side, we extracted commits that touched exactly one file and whose modified lines were confined to a single function. For each candidate commit, we inspected up to five subsequent commits that touched the same function. If any of these later messages contained problematic keywords (e.g., “revert,” “rollback,” “fix bug,” “regression”), the candidate was discarded to avoid false-clean labels—since such signals in subsequent edits suggest that the earlier commit may have introduced issues—an idea conceptually similar to SZZ methodology. Surviving candidates were further required to be modified at least once after their introduction, which excluded dead or unused code. We then formed an evidence bundle—commit message, function_before, and function_after— and queried GPT-4o three independent times under a dedicated clean-label rubric. The rubric classified each commit as: (A) Bug fix (DefectiveClean), (B) Improvement/optimization/feature change (CleanClean), or (C) Ambiguous/irrelevant commits. Categories (A) and (B) were retained as clean modifications, while any case with a (C) vote was filtered out. The sample was kept only when all three votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B).
Table 1 summarizes the filtering outcomes. The defective pool was reduced from 6,257 to 3,164—about half of revert-anchored commits were discarded, mostly because the reasons for the reverts were not sufficiently specified. Among the retained cases, 1,270 were classified as DefectiveDefective and 1,795 as CleanDefective. By contrast, the clean pool was only modestly reduced (11,807 to 10,268), with 6,524 DefectiveClean and 3,744 CleanClean cases retained, all meeting the unanimous voting criterion. The exact GPT prompts and rubric are available in the provided artifact.
| Category | Before GPT Filtering | After GPT Filtering |
| Defective total | 6,257 | 3,164 |
| DefectiveDefective | – | 1,270 |
| CleanDefective | – | 1,795 |
| Clean total | 11,807 | 10,268 |
| DefectiveClean | – | 6,524 |
| CleanClean | – | 3,744 |
Clean selection was intentionally conservative: a function had to be revised at least once after the candidate commit, and no problematic keywords could appear in the next five commits touching that function. Unlike prior JIT corpora (e.g., SZZ-derived (Kim et al., 2008; Kamei et al., 2016)) or static snapshot datasets (e.g., Devign (Zhou et al., 2019), Big-Vul (Fan et al., 2020), Defects4J (Just et al., 2014)) that define “clean” at a snapshot or immediate post-fix state, our per-function, post-hoc history check aimed to reduce false-clean labels arising from late-emerging faults or follow-up fixes.
After assembling the labels, each repository was split time-respectively (80%/10%/10% into train/validation/test by commit timestamp) to avoid temporal leakage—training on commits occurring after the test set would otherwise expose future information. Class balance (defective/clean) was preserved within each split, and splitting was performed per repository to ensure that every project was represented across partitions, preventing cases where a project might be entirely absent from validation or test.The per-project partitions were then concatenated to form the final splits.
To characterize input length, we tokenized function_after using each of the PLMs employed in our experiments (CodeBERT, UniXcoder, CodeT5+). As shown in Figure 3, the cumulative distribution indicates that although the exact counts vary slightly by model, around 55–65% of functions exceed the 512-token limit, meaning the entire function cannot be fed to the model without truncation. With UniXcoder’s 1,024-token limit, roughly 70% of functions fit in full. This distribution motivates a comparison between encoding strategies that provide only the edited lines versus those that supplying the full function at the risk of truncation. Although the dataset is derived from commit history, the intended inference unit is a function-level modification, which can be evaluated pre-commit (e.g., in an Integrated Development Environment (IDE) or local CI hooks) as well as at review/merge time. To assess label reliability, two authors independently audited a stratified sample of 100 defective commits under a shared rubric and adjudicated disagreements.
3.2. PLM Evaluation Protocol
Our goal is to assess whether pre-trained language models (PLMs) can effectively leverage code-change information for defect prediction. To this end, we designed a two-stage evaluation. Stage 1 compares alternative input encodings that vary in how much context and change signal are provided. Stage 2 introduces counterfactual perturbations that deliberately distort change signals, checking whether models degrade as expected when edit semantics are disrupted. Together, these stages provide a structured protocol for evaluating whether PLMs rely on genuine semantic cues rather than superficial patterns.
In Stage 1, we compared five alternative input encodings, each tokenized using the model’s native tokenizer before being fed into the PLM encoder.
After-only: The function_after, without any explicit indication of where edits occurred. Serves as a classical SDP-style baseline.
After+Change markers: The function_after, augmented with <CHG> tags on each modified line so the model can directly identify edited regions.
Before+After: Both pre- and post-modification functions concatenated, prefixed with markers (“[Before]: …[After]: …”). To balance length, half of the token budget was allocated to each side, ensuring that the after-code was not entirely truncated by a long before-code.
Diff with tags: Only the changed lines and their local context, with <ADD> and <DEL> tags prepended to additions and deletions respectively, following the representation of DeepJIT (Hoang et al., 2019).
AddedDeleted blocks: Inspired by JiT-Fine (Ni et al., 2022), we supplied added lines first, then deleted lines, each under explicit headers (“[ADDED LINES]: …[DELETED LINES]: …”). Unlike JiT-Fine, commit messages were excluded for fairness.
The first three variants serve as heuristic baselines, representing straightforward ways of exposing code modifications to PLMs without relying on prior proposals. These five encodings let us quantify the trade-off between context and change signal under realistic token budgets and identify the most effective representation for defect prediction.
In Stage 2, we applied controlled perturbations to test whether models behave consistently when change signals are deliberately altered. Each perturbation targeted the representation of change information while leaving the overall code form largely intact (Ribeiro et al., 2018; Jia and Liang, 2017).
Spurious change markers: In the After+Change markers format, the number of modified lines is first computed, then the same number of <CHG> tags are randomly inserted at arbitrary locations, injecting meaningless signals during both training and testing.
Swapped snapshots: Under the Before+After format, models are trained identically, but at test time the positions of function_before and function_after are swapped with 50% probability (e.g., “[Before]: function_after [After]: function_before”).
Reversed diff tags: Under the Diff with tags format, models are trained identically, but tested with the <ADD> and <DEL> tags inverted.
Swapped added/deleted blocks: In the AddedDeleted blocks format, the two sections are randomly swapped with 50% probability, so that “[ADDED LINES]” contains deletions and “[DELETED LINES]” contains additions during both training and testing.
These counterfactual variants are designed to reduce model performance: if a PLM truly understands the relationship between code before and after—that is, the semantics of the modification—then disrupting this information should lead to a drop in performance compared to the corresponding Stage 1 results. Conversely, if performance remains largely unchanged, it suggests that the model has not learned meaningful change semantics and is instead relying on superficial cues from the overall code context.
4. Experimental Setup
4.1. Research Questions
We try to answer following research questions:
RQ1: How accurate are the labels in the ReDef JIT-SDP dataset?
RQ2: Which input encodings yield the best PLM performance?
RQ3: Does PLM performance degrade under counterfactual perturbations of change representations?
4.2. Models and Baselines
To evaluate how PLMs capture the semantics of code changes, we experimented with the following pre-trained models. CodeT5+ and CodeBERT were used with their default maximum input limit of 512 tokens. UniXcoder uniquely supports longer contexts, so we evaluated it under both 512- and 1,024-token limit settings to test performance under constrained and full-capacity conditions.
Pre-trained Language Models.
CodeBERT. (Feng et al., 2020) We included CodeBERT, a 125M parameter RoBERTa-based encoder pre-trained for code understanding. Despite being one of the earlier PLMs, it has consistently shown strong and robust performance in defect prediction and classification tasks, making it an essential baseline for fair comparison (Xiao et al., 2023; Dahou et al., 2025).
CodeT5+. (Wang et al., 2023b) We used the 220M encoder-only variant, as detection tasks do not require generation capabilities. CodeT5+ has achieved state-of-the-art results across a wide range of software engineering tasks (Niu et al., 2023; Gao et al., 2024), making it a competitive and modern baseline.
- (3)
In addition to these PLMs, we implemented input encodings inspired by DeepJIT (diff with <ADD>/<DEL> tags) and JIT-Fine (addeddeleted block format). These were not used as standalone baselines, but rather as representation strategies integrated into our Stage 1 evaluation protocol.
4.3. Evaluation Metrics
We evaluate model performance using five standard classification metrics: Accuracy, Precision, Recall, F1 (F1-score), and PR-AUC. Among these, F1 is adopted as the primary metric, as it balances precision and recall and is widely used for imbalanced datasets such as JIT-SDP (Hoang et al., 2019; Kamei et al., 2012; Hoang et al., 2020).
Accuracy measures the overall proportion of correctly classified modifications, but can be misleading under severe class imbalance.
Precision quantifies the fraction of predicted defective modifications that are truly defective, while Recall measures the fraction of actual defective modifications that are successfully identified.
F1 is the harmonic mean of precision and recall, and is widely regarded as the most informative single measure in imbalanced classification settings.
PR-AUC is the area under the Precision–Recall curve, which provides a threshold-independent view of the trade-off between precision and recall across all operating points.
4.4. Implementation Details
For each PLM–encoding configuration, we fine-tuned the model ten times with different random seeds (12345–12354) and report the mean and standard deviation to reduce stochastic variance. All models were trained under consistent hyperparameters: a learning rate of , batch size 8, and the AdamW optimizer with decoupled weight decay (Loshchilov and Hutter, 2017). Gradient accumulation was applied to stabilize updates, and a linear learning-rate scheduler with a 10% warm-up ratio was used to avoid abrupt changes in the early stages. Each run lasted for 6 epochs.
The best checkpoint was selected based on validation F1 (our primary metric) and subsequently used for test reporting. To address class imbalance, we applied the standard weighting scheme, assigning the positive class a weight of (King and Zeng, 2001; Lin et al., 2017).
All experiments were conducted on Tesla V100-PCIE-32GB GPUs in a shared academic cluster. Due to CPU bottlenecks and scheduling constraints, full GPU capacity was not always utilized. Training time was consistent across encoding strategies since only the input format varied; for reference, a single CodeBERT run at the 512-token limit required approximately 30 minutes on one GPU.
Statistical Analysis. For RQ2, we conducted a two-way repeated-measures ANOVA with within-subject factors model (4) and encoding (5), treating the 10 seeds as subjects. We report partial and Cohen’s as effect sizes (Lakens, 2013). Post-hoc comparisons used paired -tests with Holm step-down correction (Holm, 1979) applied per family: Encoding main-effect pairs on subject-level marginal means (collapsed over models), and all encoding pairs within each model (simple effects).
For RQ3, each perturbation was paired with its corresponding original encoding and evaluated across the same 10 seeds. We tested the directional hypothesis that perturbations degrade performance using one-sided paired -tests (), complemented by Wilcoxon signed-rank tests (Wilcoxon, 1945). We also report paired effect size as Cohen’s and non-parametric percentile 95% bootstrap confidence intervals (CIs) for the mean difference using 10,000 resamples (Tibshirani and Efron, 1993). -values were adjusted for multiple comparisons using Holm’s method within each set of tests (e.g., across the five perturbations).
Following common conventions, we consider results statistically significant at . Effect sizes are interpreted using standard thresholds: partial (small), (medium), (large); Cohen’s (small), (medium), (large) (Cohen, 2013).
5. Experimental results
RQ1: How accurate are the labels in the ReDef JIT-SDP dataset?
We manually audited 100 randomly sampled defective functions. Following recent practices in vulnerability dataset curation (Chen et al., 2023; Ding et al., 2024), two authors independently inspected the commit evidence—reverted commit message, function_before, function_after, and revert commit message—and labeled each case as Correct, Ambiguous, or Wrong. After adjudication, 92 cases were classified as Correct, 5 as Ambiguous, and 3 as Wrong, yielding a positive-label precision of 0.92 as shown in Table 2. Compared to prior reports of 14–50% noise in SDP datasets (Herbold et al., 2022; Afric et al., 2023), our 8% mislabel rate is substantially lower.
| Correct | Ambiguous | Wrong | |
| Reviewer 1 | 94 | 5 | 1 |
| Reviewer 2 | 83 | 14 | 3 |
| Adjudicated | 92 | 5 | 3 |
Ambiguities initially accounted for 30% of cases, typically arising when revert messages cited only a hash without any rationale, forcing annotators to infer defect-relatedness. We revised the rubric to abstain unless defect links were explicit, which reduced ambiguous cases to 5% in the final adjudicated set. The three Wrong cases followed non-defect patterns (e.g., mistaken commit ordering, temporary mitigations, or structural refactoring).
Overall, these results suggest that our revert-anchored, history-screened pipeline yields comparatively reliable defect labels.
RQ2: Which input encodings yield the best PLM performance?
Table 3 summarizes the performance of all PLM–encoding combinations. Overall, Diff w/ tags and AddedDeleted consistently achieved the highest F1 across models, whereas After+Markers and Before+After performed worst. In contrast, these latter two encodings often ranked highest in terms of accuracy. This discrepancy is explained by class imbalance: ReDef contains over three times more clean (negative) than defective (positive) modifications, so encodings biased toward the majority class can appear more accurate while sacrificing F1. To compare models more rigorously beyond such raw patterns, we applied statistical tests as follows.
| Model | Encoding | F1 | PR-AUC | Accuracy | Precision | Recall |
| CodeBERT | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| CodeT5+ | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| UniXcoder-512 | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| UniXcoder-1,024 | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted |
We conducted a two-way repeated-measures ANOVA with model (4 PLMs) and encoding (5 variants) as within-subject factors (Table 4). Results showed significant main effects of both model () and encoding (), both indicating large effect sizes. No interaction was found (), suggesting that the relative ranking of encodings was consistent across PLMs. Post-hoc pairwise tests with Holm correction revealed that Diff w/ tags and AddedDeleted significantly outperformed After+Markers and Before+After (all ), while performing comparably to each other. After-only showed intermediate performance, significantly better than some weaker encodings but worse than the top two in most cases. The pattern was consistent for CodeBERT, CodeT5+, and UniXcoder-512, whereas for UniXcoder-1,024 the differences diminished and were not statistically significant.
| Effect | df_num | df_den | ||||
| Model | 9.96 | 3 | 27 | 0.53 | 1.05 | |
| Encoding | 13.80 | 4 | 36 | 0.61 | 1.24 | |
| ModelEncoding | 0.82 | 12 | 108 | .63 | 0.08 | 0.30 |
Across models, this advantage of change-focused encodings can be explained by two factors. First, as shown in Figure 3, only about 30–40% of functions fit within 512 tokens, so encodings that pass the entire function (After-only, After+Markers, Before+After) often truncate important edits. Second, supplying the entire function can also introduce excessive, mostly unchanged context that distracts the model from the actual modification. The weakness of Before+After illustrates both issues: two full functions must fit under the same budget, increasing truncation risk and noise. Supporting the second explanation, UniXcoder performed worse at 1,024 tokens than at 512, suggesting that longer sequences add distracting context rather than improving focus. Finally, in our setting which primarily targets code understanding, CodeBERT again emerged as the strongest PLM, in line with recent studies showing its robustness on classification and comprehension tasks (Xiao et al., 2023; Dahou et al., 2025).
RQ3: Does PLM performance degrade under counterfactual perturbations of change representations? As shown in Table 5, across the four perturbations—Spurious change markers, Swapped snapshots, Reversed diff tags, and Swapped added/deleted blocks—performance remained effectively unchanged. Even when before/after alignment or polarity direction was disrupted, F1 differences stayed below 1%. This stability is counterintuitive: if PLMs were truly leveraging semantic signals from code changes, such distortions should have produced clear drops in performance metrics. Instead, the negligible differences suggest that the models are not semantically sensitive to modification content, but rather exploit surface-level regularities (e.g., token distributions, positional patterns) that persist under perturbation. In other words, the robustness observed here reflects insensitivity rather than genuine understanding.
Percent change is relative to the corresponding original encoding.
| Model | Condition | F1 | PR-AUC | Accuracy | Precision | Recall |
| CodeBERT | After+Markers | |||||
| Spurious CHG markers | ||||||
| +0.23% | ||||||
| CodeBERT | Before+After | |||||
| Swapped snapshots | ||||||
| +0.48% | ||||||
| CodeBERT | Diff w/ tags | |||||
| Reversed diff tags | ||||||
| CodeBERT | AddedDeleted | |||||
| Swapped added/deleted blocks | ||||||
Statistical tests (paired t-tests, Wilcoxon signed-rank tests, bootstrap with Holm correction) confirmed this robustness. Only Reversed diff tags showed statistical significance ( after correction), but the absolute effect size was negligible (, Cohen’s ). In practice, such a tiny change does not indicate genuine semantic sensitivity. Taken together, these results reveal an important limitation: current PLMs achieve stable performance not because they robustly model code-change semantics, but because they fall back on shallow cues that remain unaffected by perturbations. This highlights a fundamental gap between snapshot-level code understanding and true modification reasoning, and points to the need for edit-aware pre-training objectives or architectures explicitly designed to capture change semantics.
6. Threats to Validity
6.1. Internal Validity
Choice of Input Encoding. We evaluated five representative encodings, including formats inspired by prior JIT-SDP studies. However, given that all encodings ultimately derive from function_before and function_after, other ways of structuring or combining these views may provide richer signals of semantic change. For example, tree-structured or graph-based encodings could expose different aspects of program structure, and hybrid schemes may better capture interactions between edits and surrounding context. Exploring such alternatives remains an open avenue for future work.
Statistical Power. All statistical tests were based on 10 independent runs with different seeds. While this design controls stochastic variance, a larger number of repetitions could yield stronger statistical power and narrower confidence intervals. However, prior AI studies on code intelligence tasks rarely report extensive repetitions: many omit random seeds entirely or rely on a single run, and even the more rigorous works seldom exceed 10 repetitions due to the substantial computational cost of PLMs (Ye et al., 2023; Grishina et al., 2023). Running 30 or more trials per configuration is generally infeasible in practice. Our choice of 10 runs follows this common practice while still providing a reasonable trade-off between robustness and resource constraints.
6.2. External Validity
Project Selection Bias. Our dataset was drawn from large, long-lived infrastructure projects (e.g., Linux, FFmpeg, Chromium). These may differ from smaller applications or libraries in review culture, test maturity, and revert policies. We partially mitigated this risk by including diverse domains and reporting a leave-one-project-out analysis, but broader coverage is needed. In particular, industrial or proprietary projects could exhibit markedly different defect patterns, and including such settings would provide stronger external validity.
Cross-Language Applicability. This study focused solely on C/C++ projects. Substantial syntactic and semantic differences across programming languages may limit the applicability of our findings. Extending evaluation to multi-language datasets is a key direction for future work. Such efforts should examine whether our observations hold consistently across different language families.
7. Conclusion
This work introduced ReDef, a high-confidence JIT-SDP dataset of function-level modifications with both defective and clean cases: 3,164 defective and 10,268 clean function-level modifications, with manual audits confirming 92% label precision—substantially higher than the prior JIT-SDP corpora. Our curation pipeline combines revert anchoring, longitudinal history screening, and GPT-assisted triage with multiple votes and audits to reduce label noise. Unlike conventional SZZ-based resources which infer bug-inducing commits post hoc and often lack clean counterparts, ReDef offers the community a reliable benchmark for studying change-aware models, enabling reproducible comparison and facilitating exploration of edit-sensitive learning objectives.
Using ReDef, we fine-tuned three strong PLMs—CodeBERT, CodeT5+, and UniXcoder—under five input encodings. Compact, change-focused encodings and tagged diffs achieved the best results, whereas whole-function inputs often diluted the signal and suffered from truncation under token limits. Notably, UniXcoder performed worse at a 1,024-token budget than at 512, showing that more context does not necessarily help when change information becomes obscured.
Counterfactual perturbations of change representations yielded almost no performance degradation. Even when we swapped the order of deleted and added blocks or mislabeled unchanged lines as modified, F1 remained nearly identical. If PLMs had truly internalized the semantics of edits, such distortions should have impaired performance. This outcome exposes a critical blind spot in how current PLMs are evaluated on code modifications: what appears to be robustness is, in fact, an artifact of surface-level cue reliance. Our findings caution against interpreting strong benchmark scores as evidence of genuine change understanding.
Threats to validity remain. We examined only five encoding strategies, but alternative ways of presenting function_before and function_after may provide richer semantic signals. Our dataset focuses on large C/C++ projects, leaving open questions about smaller repositories and other languages. Finally, although we repeated all experiments with ten random seeds, stronger statistical power could be obtained with more runs.
Future work should probe why PLMs lag in modification understanding compared to single-snapshot tasks such as vulnerability detection. Promising directions include pre-training objectives that explicitly expose models to edits and diffs, as well as richer change representations or architectures that make modification semantics more accessible to PLMs. Beyond PLMs, evaluating decoder-based code LLMs (e.g., CodeLlama, StarCoder) and API models (e.g., GPT-4o) under the same protocol will test whether larger generative models can close this gap.
We hope that ReDef and our evaluation protocol will serve as a foundation for the next generation of change-aware, semantics-sensitive defect prediction models.
Data Availability
An anonymous link containing the dataset, construction scripts, and experimental code is publicly available for reproducibility:https://figshare.com/s/4f202bc0921e26b41dc2.