### v2

# ReDef: Do Code Language Models Truly Understand Code Changes for Just-in-Time Software Defect Prediction?

Abstract.
Just-in-Time software defect prediction (JIT-SDP) plays a critical role in prioritizing risky code changes during code review and continuous integration. However, existing datasets often suffer from noisy labels and low precision in identifying bug-inducing commits. To address this, we present ReDef (Revert-based Defect dataset), a high-confidence benchmark of function-level modifications curated from 22 large-scale C/C++ projects. Defective cases are anchored by revert commits, while clean cases are validated through post-hoc history checks. Ambiguous instances are conservatively filtered out via a GPT-assisted triage process involving multiple votes and audits. This pipeline yields 3,164 defective and 10,268 clean modifications, offering substantially more reliable labels than prior resources. Beyond dataset construction, we provide a systematic evaluation of how Code Language Models (CLMs)—specifically CodeBERT, CodeT5+, UniXcoder, and Qwen2.5—reason about code modifications. We first investigate which input encodings most effectively expose change information under five different strategies. We then design four counterfactual perturbation strategies (e.g., swapping added/deleted blocks, inverting diff polarity) to serve as diagnostic probes. We posit that if models genuinely capture change semantics, such distortions should lead to a clear decline in predictive performance. Our results show that compact diff-style encodings consistently outperform whole-function formats across all CLMs, supported by rigorous statistical confirmation. However, under counterfactual tests, performance remains effectively stable, revealing that what appears to be robustness in fact reflects a reliance on superficial cues rather than true semantic understanding. These findings indicate that, at least in code-change understanding tasks, current CLMs remain limited in their ability to genuinely comprehend the relational dynamics of code modifications.
1. Introduction
Software defect prediction (SDP) aims to estimate whether a given piece of code is likely to contain defects so that limited engineering resources—testing, review time, and Continuous Integration (CI) capacity—can be prioritized where they matter most (Malhotra, 2015). Unlike defect detection or localization, which seeks to identify concrete bug instances, SDP is a risk estimation problem used to triage changes and guide quality assurance activities. Effective SDP reduces downstream failures and lowers maintenance costs by surfacing risky code early enough for developers to take action within their normal workflow—for example, during code review or before a merge in CI.
Classical SDP typically models static snapshots of code or aggregates hand-crafted process metrics—such as churn, complexity, ownership, and past defects—at the file, module, or function level (Wahono, 2015). While such predictors can correlate with risk, they fail to capture how the code is evolving at the moment of change. Just-in-Time (JIT) SDP addresses this limitation by reasoning over code modifications at commit time, aligning closely with practical review and CI processes that scrutinize the changed lines (Kamei et al., 2012; Kim et al., 2008). However, most existing JIT-SDP datasets rely on SZZ(Śliwerski–Zimmermann–Zeller)-style linkages from bug-fixing commits back to bug-introducing commits (Śliwerski et al., 2005). SZZ attributes the introduction of a defect to the last modifier of the lines changed by a later fix—an assumption that is often violated by routine edits such as renaming or refactoring. This approach introduces substantial noise: it can misidentify the true introduction points, propagate commit-level label errors, and mismatch the actual defective lines, ultimately obscuring evaluation quality (Herbold et al., 2022; Lyu et al., 2024). To mitigate the noise inherent in SZZ-based labeling, we turn to a more explicit and high-confidence signal: revert commits.
Reverts are a routine part of disciplined software development. Developers may revert commits for various reasons—such as regressions, incompatibilities, performance degradations, policy or formatting violations, or simply because another change has superseded the patch (Yan et al., 2019). A revert creates an explicit, time-stamped link to the earlier change. When the revert message cites a failure or fix, it provides strong evidence that the prior modification was defective. We therefore use revert commits as anchors to harvest high-confidence positive examples, and subsequently filter out cases unrelated to defective modifications.
With a more reliable labeling strategy in place, the next question is how modern models can be effectively leveraged. Existing JIT-SDP research has largely depended on compact engineered features with classical machine learning (ML) and, more recently, deep learning (DL) models. However, these approaches primarily capture surface-level correlations—such as churn, ownership, and developer activity—rather than the true semantics of code changes (Kamei et al., 2012, 2016; Hoang et al., 2019). In other words, they estimate probabilities from shallow cues instead of reasoning about what the code actually does.
By contrast, recent Code Language Models (CLMs)—ranging from specialized encoder-based architectures to large-scale generative decoders (LLMs)—have demonstrated an unprecedented ability to capture the deep semantic meaning of code, moving beyond mere statistical correlates. This advancement has led to state-of-the-art performance across diverse software engineering tasks, including code search, summarization, repair, and vulnerability detection (Guo et al., 2022; Xia et al., 2023; Ni et al., 2023; Hui et al., 2024). This success raises a fundamental question for JIT-SDP: if CLMs can effectively understand static code snapshots, can they also reason about code modifications (i.e., the causal relationship between the before and after states)? Given that CLMs, regardless of their specific architecture, have consistently shown superior effectiveness in code understanding, we focus our study on evaluating these models as the primary candidates for a systematic evaluation of change-level defect prediction (Zhang et al., 2025b; Sun et al., 2023; Zhang et al., 2025a).
To answer this question, we first construct a high-confidence, change-centric benchmark. This high label precision is essential for a rigorous analysis, as it ensures that any observed model limitations stem from a lack of semantic understanding rather than noise in the ground truth. Next, in Stage 1, we compare alternative ways of presenting modification information to CLMs, asking which input encodings yield the most effective performance. For this, we fine-tune four widely used CLMs—CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), UniXcoder (Guo et al., 2022), and Qwen2.5 (Hui et al., 2024)—as backbones. However, strong results under these encodings alone do not guarantee genuine semantic reasoning, as models may still rely on superficial cues. Therefore, in Stage 2, we deliberately perturb the representation of code changes—e.g., by swapping added and deleted code blocks, inverting edit polarity, or injecting spurious markers—to test whether CLMs degrade when true semantic information is distorted. By probing how models behave under such distortions, our protocol not only measures raw accuracy but also uncovers a blind spot in current CLM research: apparent robustness may simply reflect reliance on shallow cues rather than genuine semantic reasoning about code changes.
Our key contributions are as follows:
High-Confidence JIT-SDP Dataset (ReDef): We curate ReDef (Revert-based Defect dataset), a corpus of single-function changes harvested via revert anchors and refined through GPT-assisted filtering. The dataset is further validated through manual sampling to ensure benchmark reliability, and its 10K+ (3,164 defective, 10,268 clean) scale makes it well-suited for fine-tuning CLMs.
Encoding Strategies for CLMs: We conduct the first systematic comparative study of five encoding strategies for representing code modifications as CLM inputs, providing clear empirical evidence of how input representations affect change-centric defect prediction.
Counterfactual Robustness Analysis: We design counterfactual perturbation tests to probe whether CLMs truly reason about code changes beyond surface cues, revealing limitations of current models and offering guidance for future research.
Together, these contributions provide both a reliable benchmark and empirical insights into how CLMs handle code changes, informing future defect prediction research and practice. The remainder of this paper is organized as follows. Section 2 reviews background and related work. Section 3 describes our dataset construction and two-stage evaluation protocol, encompassing both input encoding strategies and counterfactual perturbation methods. Section 4 outlines the experimental setup, and Section 5 reports results. Section 6 discusses threats to validity, and Section 7 concludes.
2. Background and Related Work
This section reviews prior research on software defect prediction and the challenges of dataset reliability. Section 2.1 briefly outlines classical Software Defect Prediction (SDP), which focuses on identifying defect-prone modules from static snapshots. Section 2.2 covers the evolution of Just-in-Time software defect prediction (JIT-SDP), focusing on the transition from hand-crafted metrics to deep learning and recent Code Language Models (CLMs). Finally, Section 2.3 discusses the critical limitations of existing JIT-SDP datasets, specifically the noise and reliability issues inherent in SZZ-based labeling, which motivates our construction of the ReDef benchmark.
2.1. Artifact-level Software Defect Prediction (SDP)
Classical SDP focuses on predicting defect-prone artifacts (e.g., files or modules) from static snapshots. Early approaches relied on hand-crafted product metrics (e.g., size, complexity, coupling) and process metrics (e.g., churn, developer ownership, past defects) (Malhotra, 2015; Bell et al., 2013; Huda et al., 2017) fed into traditional machine learning models like SVMs and random forests (Elish and Elish, 2008; Balaram and Vasundra, 2022; Wang et al., 2012; Khoshgoftaar and Seliya, 2002). More recent work has shifted toward learned representations using deep generative models or transformer-based encoders to capture richer features (Wang et al., 2018; Omri and Sinz, 2020). While these methods have evolved to consider various practical evaluation settings (Ostrand and Weyuker, 2007; Tosun et al., 2009), snapshot-based prediction remains limited as it cannot capture the semantics of specific code changes and often blurs defective regions at a coarse granularity. This lack of change-level precision necessitates Just-in-Time (JIT) approaches that provide actionable insights at the time of commit.
2.2. Just-in-Time Software Defect Prediction (JIT-SDP)
JIT-SDP shifts the focus from static artifacts to change-time risk estimation, scoring each commit or patch so that reviewers and CI gates can provide immediate feedback. The concept originates from Kim et al., who first classified individual changes as clean or buggy (Kim et al., 2008). Subsequent work engineered compact commit-level features—such as churn, diffusion, prior defect density, developer experience, recency, and message cues—and applied machine learning techniques (Kamei et al., 2012; Shihab et al., 2012; Kamei et al., 2016). Later studies incorporated commit messages and patch text, while neural models began encoding diffs directly. For example, Yang et al. proposed Deeper, a Deep Belief Network (DBN)-based model that outperformed logistic regression (Yang et al., 2015), and Hoang et al. introduced DeepJIT, which integrates commit messages and code metrics (Hoang et al., 2019). CC2Vec further embedded code changes hierarchically to capture structural and sequential information (Hoang et al., 2020).
More recently, researchers have applied CLMs to JIT-SDP. Guo et al. systematically compared six transformer-based backbones (RoBERTa, CodeBERT, BART, PLBART, GPT2, and CodeGPT) and showed that CLM-based variants consistently outperform traditional baselines such as DeepJIT and CC2Vec (Guo et al., 2023). Building on this, Abu Talib et al. demonstrated that parameter-efficient fine-tuning (PEFT) techniques such as LoRA can further boost JIT-SDP performance on ApacheJIT (Abu Talib et al., 2024), while Lin et al. introduced CCT5—a CodeT5-initialized model pre-trained on CodeChangeNet—that achieves state-of-the-art results on code-change tasks including JIT-SDP (Lin et al., 2023). Jiang et al. extended CodeBERT with a Replaced Message Identification objective, resulting in BiCC-BERT and JIT-BiCC, which improved F1-scores by over 10% (Jiang et al., 2025). Together, these studies highlight the promise of CLMs for JIT-SDP, appearing to possess an advanced capability to comprehend code modifications. However, most existing research has focused on compact encoder-based or encoder-decoder models, leaving the potential of contemporary large-scale decoder-based models—such as the Qwen (Yang et al., 2025) or Llama families (Grattafiori et al., 2024)—largely underexplored in this domain. Furthermore, it remains unclear whether these performance gains reflect genuine semantic understanding or a reliance on superficial textual patterns, underscoring the need for a systematic evaluation of model behavior.
2.3. Labeling Reliability and SZZ Limitations
A precise and high-confidence dataset is a fundamental prerequisite for evaluating whether Code Language Models truly understand the semantics of code changes. Despite the technical progress in neural modeling described in Section 2.2, a critical bottleneck in JIT-SDP remains the reliability of dataset labeling. Most existing JIT-SDP datasets rely on SZZ-based labeling, which identifies bug-introducing commits by tracing the origins of the code lines modified in bug-fixing commits (Śliwerski et al., 2005). However, this approach introduces substantial noise that propagates labeling errors and obscures evaluation quality. Herbold et al. showed that SZZ-based labeling is highly error-prone, with only about half of the commits identified as bug-inducing being correct (Herbold et al., 2022). Even after additional refinement using a RoBERTa-based issue classifier, Afric et al. reported that datasets still contained, on average, 14.3% mislabeled instances (Afric et al., 2023). Lyu et al. conducted a large-scale evaluation on over 76K developer-labeled Linux kernel commits and showed that SZZ variants suffer from a recall degradation of 13.8%, with 17.47% of bug-fixing commits being ghost commits (Lyu et al., 2024).
To mitigate these errors, several SZZ variants have been proposed (Fan et al., 2019). Early improvements like Annotation Graph SZZ (AG-SZZ) attempted to filter non-semantic changes (e.g., whitespace and comments) (Kim et al., 2006). Subsequent variants addressed meta-changes (MA-SZZ) (Da Costa et al., 2016) and automated refactoring detection (RA-SZZ) (Neto et al., 2018) to prevent incorrect blame assignment. More recent efforts include TC-SZZ, which iteratively traces deeper histories to recover missed commits (Lyu et al., 2024), and Code Change Tactics (CAT), which disentangles refactoring-propagated changes from actual defect-inducing edits (Niu et al., 2025). Yet, even with these sophisticated refinements, automated labeling still struggles with the inherent ambiguity of commit histories.
Fundamentally, pinpointing the root cause of a defect requires deep domain expertise; without explicit documentation in commit messages, it is notoriously difficult for external researchers—or automated tools—to accurately assign blame to specific code modifications. Furthermore, even manual inspection by non-developers remains subjective and error-prone, as identifying a bug’s origin from raw diffs alone lacks necessary semantic grounding. These persistent reliability gaps underscore the need for more conservative and evidence-based labeling strategies that rely on explicit developer feedback rather than speculative history tracking. In this paper, we address this need by introducing the ReDef benchmark, which prioritizes high-confidence signals to provide a more robust environment for assessing the true reasoning capabilities of various Code Language Models, spanning encoder-based models (CodeBERT, UniXcoder, and the encoder of CodeT5+) and a large-scale decoder-based model (Qwen2.5).
3. Methodology
Our methodology comprises two main pillars. (A) ReDef Construction: We curate defective and clean code modifications from large-scale C/C++ projects, leveraging revert signals, post-hoc history screening, and expert audits to ensure high label reliability. (B) Code Language Models Evaluation Protocol: We design a two-stage framework to assess how various Code Language Models (CLM) interpret code modifications. Figure 1 summarizes the overall approach.
3.1. ReDef Construction
We curated ReDef, a high-confidence JIT-SDP corpus from 22 long-lived, high-activity C/C++ repositories. Projects were selected for sustained commit volume, explicit revert workflows, and domain diversity to reduce project-specific bias. The curation yielded 3,164 defective and 10,268 clean modifications. Figure 2 reports per-project counts.
Defective Modification. We extracted commits that touched exactly one file and whose modified lines were confined to a single function. This design ensures reliable causal attribution to a specific function—a choice also adopted in recent datasets such as PrimeVul (Ding et al., 2024). From this single-function pool, we identified defective modifications by collecting revert commits whose messages (i) contained the keyword “revert” and (ii) explicitly referenced the hash of an earlier commit. Since such reverts typically indicate that the referenced commit introduced a problem, the targeted commits are highly likely to represent defective modifications.
For each referenced hash, we retrieved its associated commit (the target of the revert) and collected four artifacts for analysis: the reverted commit message, the function_before and function_after from that commit, and the revert commit message itself. Together these formed an evidence bundle and were submitted to GPT-4o three independent times under a fixed rubric to remove ambiguous or non-defect reverts (e.g., formatting changes, policy reversions, or hash-only messages without rationale). The rubric distinguished each revert as: (A) Failed/incomplete fix (DefectiveDefective), (B) Bug introduction (CleanDefective), or (C) Other (non-defect). Categories (A) and (B) were merged into the defective class, while (C) served as a filter to discard non-defect cases. To ensure high precision, a candidate was retained only if all three GPT-4o votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B). Any sample with one or more (C) votes (e.g., A/A/C) was conservatively discarded.
Across the 22 subject projects, we identified approximately 0.43M reverts out of 3.37M fix-related commits, indicating that reverts account for roughly 13% of the total defect pool. This proportion further decreases after applying the single-function constraint and GPT-4o triage. While this design inevitably limits coverage, it reflects a deliberate trade-off to prioritize label precision and mitigate the “falsely blaming” issue inherent in traditional SZZ-based heuristics. In contrast to heuristic lineage tracking, revert commits provide explicit causal links between bug fixes and their corresponding bug-inducing changes, and our filtering pipeline removes tangled or ambiguous cases. As a result, ReDef complements existing larger but noisier datasets by offering a controlled, high-confidence benchmark for evaluating the semantic reasoning capabilities of code models.
Clean Modification. To assemble the clean stratum, we first identified all defective modifications and then targeted clean modification commits from a similar time period in the same repository, thereby mitigating both project-specific and period-specific bias. As with the defective side, we extracted commits that touched exactly one file and whose modified lines were confined to a single function. For each candidate commit, we inspected up to five subsequent commits that touched the same function. If any of these later messages contained problematic keywords (e.g., “revert,” “rollback,” “fix bug,” “regression”), the candidate was discarded to avoid false-clean labels—since such signals in subsequent edits suggest that the earlier commit may have introduced issues—an idea conceptually similar to SZZ methodology. Surviving candidates were further required to be modified at least once after their introduction, which excluded dead or unused code. We then formed an evidence bundle—commit message, function_before, and function_after— and queried GPT-4o three independent times under a dedicated clean-label rubric. The rubric classified each commit as: (A) Bug fix (DefectiveClean), (B) Improvement/optimization/feature change (CleanClean), or (C) Ambiguous/irrelevant commits. Categories (A) and (B) were retained as clean modifications, while any case with a (C) vote was filtered out. The sample was kept only when all three votes unanimously fell in {A,B} (e.g., A/A/A or A/B/B).
Table 1 summarizes the filtering outcomes. The defective pool was reduced from 6,257 to 3,164—about half of revert-anchored commits were discarded, mostly because the reasons for the reverts were not sufficiently specified. Among the retained cases, 1,270 were classified as DefectiveDefective and 1,795 as CleanDefective. By contrast, the clean pool was only modestly reduced (11,807 to 10,268), with 6,524 DefectiveClean and 3,744 CleanClean cases retained, all meeting the unanimous voting criterion. The exact GPT prompts and rubric are available in the provided artifact.
| Category | Before GPT Filtering | After GPT Filtering |
| Defective total | 6,257 | 3,164 |
| DefectiveDefective | – | 1,270 |
| CleanDefective | – | 1,795 |
| Clean total | 11,807 | 10,268 |
| DefectiveClean | – | 6,524 |
| CleanClean | – | 3,744 |
Clean selection was intentionally conservative: a function had to be revised at least once after the candidate commit, and no problematic keywords could appear in the next five commits touching that function. Unlike prior JIT corpora (e.g., SZZ-derived (Kim et al., 2008; Kamei et al., 2016)) or static snapshot datasets (e.g., Devign (Zhou et al., 2019), Big-Vul (Fan et al., 2020), Defects4J (Just et al., 2014)) that define “clean” at a snapshot or immediate post-fix state, our per-function, post-hoc history check aimed to reduce false-clean labels arising from late-emerging faults or follow-up fixes.
Data Partitioning and Statistics. After assembling the labels, each repository was split time-respectively (80%/10%/10% into train/validation/test by commit timestamp) to avoid temporal leakage—training on commits occurring after the test set would otherwise expose future information. Class balance (defective/clean) was preserved within each split, and splitting was performed per repository to ensure that every project was represented across partitions, preventing cases where a project might be entirely absent from validation or test. The per-project partitions were then concatenated to form the final splits. Table 2 summarizes the scale of the curated modifications. On average, the studied functions consist of 94.52 Lines of Code (LOC), with modifications typically affecting 22.82% of the code. The gradual increase in average LOC from training to test sets reflects the natural growth of project complexity over time, as our chronological split captures the inherent evolution of software projects. Conversely, the slight decrease in the percentage of changed code suggests that modifications become more localized as projects mature and stabilize.
| Split | Count | Avg. LOC | % Changed |
| Train | 10,744 | 92.12 | 23.30% |
| Valid | 1,344 | 102.29 | 22.01% |
| Test | 1,344 | 105.98 | 19.81% |
| Overall | 13,432 | 94.52 | 22.82% |
To characterize input length, we tokenized function_after using each of the CLMs employed in our experiments (CodeBERT, UniXcoder, CodeT5+, Qwen2.5). As shown in Figure 3, the cumulative distribution indicates that although the exact counts vary slightly by model, around 48–65% of functions exceed the 512-token limit, meaning the entire function cannot be fed to the model without truncation. With Qwen2.5’s 1,024-token limit, roughly 77% of functions fit in full. This distribution motivates a comparison between encoding strategies that provide only the edited lines versus those that supplying the full function at the risk of truncation. Although the dataset is derived from commit history, the intended inference unit is a function-level modification, which can be evaluated pre-commit (e.g., in an Integrated Development Environment (IDE) or local CI hooks) as well as at review/merge time. To assess label reliability, two authors independently audited a stratified sample of 100 defective commits under a shared rubric and adjudicated disagreements.
3.2. Code Language Models Evaluation Protocol
Our goal is to assess whether Code Language Models (CLMs) can effectively leverage code-change information for defect prediction. To this end, we designed a two-stage evaluation. Stage 1 compares alternative input encodings that vary in how much context and change signal are provided. Stage 2 introduces counterfactual perturbations that deliberately distort change signals, checking whether models degrade as expected when edit semantics are disrupted. Together, these stages provide a structured protocol for evaluating whether CLMs rely on genuine semantic cues rather than superficial patterns.
In Stage 1, we compared five alternative input encodings, each tokenized using the model’s native tokenizer before being fed into the CLM encoder.
After-only: The function_after, without any explicit indication of where edits occurred. Serves as a classical SDP-style baseline.
After+Change markers: The function_after, augmented with <CHG> tags on each modified line so the model can directly identify edited regions.
Before+After: Both pre- and post-modification functions concatenated, prefixed with markers (“[Before]: …[After]: …”). To balance length, half of the token budget was allocated to each side, ensuring that the after-code was not entirely truncated by a long before-code.
Diff with tags: Only the changed lines and their local context, with <ADD> and <DEL> tags prepended to additions and deletions respectively, following the representation of DeepJIT (Hoang et al., 2019).
AddedDeleted blocks: Inspired by JiT-Fine (Ni et al., 2022), we supplied added lines first, then deleted lines, each under explicit headers (“[ADDED LINES]: …[DELETED LINES]: …”). Unlike JiT-Fine, commit messages were excluded for fairness.
The first three variants serve as heuristic baselines, representing straightforward ways of exposing code modifications to CLMs without relying on prior proposals. These five encodings let us quantify the trade-off between context and change signal under realistic token budgets and identify the most effective representation for defect prediction.
In Stage 2, we applied controlled perturbations to test whether models behave consistently when change signals are deliberately altered. Rather than simulating realistic noise, these serve as diagnostic counterfactuals to determine whether a model’s performance stems from a genuine understanding of code-change semantics (e.g., diff polarity and before/after temporal order) or from an over-reliance on superficial code features. To illustrate, consider a clean bug-fix that changes a risky null-pointer access use(p); to a safe guarded access if (p) use(p);. In the Swapped added/deleted blocks strategy, the representation is transformed as follows:
| Example: Semantic shift in a null-pointer check fix |
| Original (Clean): [DELETED LINES] use(p); [ADDED LINES] if (p) use(p); |
| A logical fix for a potential crash. |
| Perturbed (Clean): [DELETED LINES] if (p) use(p); [ADDED LINES] use(p); |
| A logical reintroduction of a bug (regression). |
If a model truly understands the semantics of a “fix,” its confidence in a “clean” label should significantly drop when presented with the perturbed version, which now logically represents a defect. A model whose predictions remain invariant to such reversals is likely relying on the presence of certain keywords (e.g., if, NULL) regardless of their roles in the change, making it unreliable for real-world deployment where distinguishing between a fix and a regression is critical.
We applied one specific perturbation to each of the four change-aware encodings introduced in Stage 1, excluding the After-only baseline as it lacks explicit change signals to disrupt. We intentionally varied the scope of these perturbations—applying them either across both training and testing or exclusively at test time—to explore two distinct diagnostic scenarios:
Spurious change markers: In the After+Change markers format, the number of modified lines is first computed, then the same number of <CHG> tags are randomly inserted at arbitrary locations. This injection of meaningless signals occurs during both training and testing.
Swapped snapshots: Under the Before+After format, models are trained on original data but perturbed only at test time by swapping the positions of function_before and function_after with a 50% probability (e.g., “[Before]: function_after [After]: function_before”).
Reversed diff tags: Under the Diff with tags format, models are trained on original data but perturbed only at test time with the <ADD> and <DEL> tags consistently inverted.
Swapped added/deleted blocks: In the AddedDeleted blocks format, the two sections are randomly swapped with a 50% probability during both training and testing, such that “[ADDED LINES]” contains deletions and “[DELETED LINES]” contains additions.
Each perturbation targeted the representation of change information while leaving the overall code form largely intact (Ribeiro et al., 2018; Jia and Liang, 2017). These counterfactual variants are designed to reduce model performance: if a CLM truly understands the relationship between code before and after—that is, the semantics of the modification—then disrupting this information should lead to a drop in performance compared to the corresponding Stage 1 results. Conversely, if performance remains largely unchanged, it suggests that the model has not learned meaningful change semantics and is instead relying on superficial cues from the overall code context.
4. Experimental Setup
4.1. Research Questions
We try to answer following research questions:
RQ1: How accurate are the labels in the ReDef JIT-SDP dataset?
RQ2: Which input encodings yield the best CLM performance?
RQ3: Does CLM performance degrade under counterfactual perturbations of change representations?
4.2. Models and Baselines
To evaluate how CLMs capture the semantics of code changes, we employed four representative models across distinct architectures. We included established encoder-based models—CodeBERT, CodeT5+, and UniXcoder—which have consistently demonstrated strong performance in JIT-SDP tasks (Jiang et al., 2025; Lin et al., 2023). Additionally, we incorporated Qwen2.5, a state-of-the-art decoder-based model. To the best of our knowledge, this study represents the first evaluation of Qwen2.5 within the JIT-SDP domain, despite its reported success in other code-related tasks.
Regarding input constraints, CodeBERT and CodeT5+ were evaluated using their default maximum limit of 512 tokens. In contrast, UniXcoder and Qwen2.5 were tested at both 512- and 1,024-token scales. For Qwen2.5 specifically, we extended the total token budget by 100 tokens (setting the limits to 612 and 1,124) to accommodate the necessary system prompt and role definitions without truncating the actual code content, thereby ensuring a fair comparison of the effective code context.
Subject Models.
CodeBERT (Feng et al., 2020) We included CodeBERT, a 125M parameter RoBERTa-based encoder pre-trained for code understanding. Despite being one of the earlier code language models, it has consistently shown strong and robust performance in defect prediction and classification tasks, making it an essential baseline for fair comparison (Xiao et al., 2023; Dahou et al., 2025).
CodeT5+ (Wang et al., 2023) We used the 220M encoder-only variant, as detection tasks do not require generation capabilities. CodeT5+ has achieved state-of-the-art results across a wide range of software engineering tasks (Niu et al., 2023; Gao et al., 2024), making it a competitive and modern baseline.
- (3)
Qwen2.5 (Hui et al., 2024) We included Qwen2.5-7B-Instruct, representing the current state-of-the-art in open-source decoder-based models. We selected the 7B variant as the standard representative of large-scale models to evaluate its reasoning capabilities over code modifications. To further distinguish between the model’s inherent zero-shot reasoning and its task-specific adaptation, we evaluated Qwen2.5 in two settings: an inference-only (zero-shot) baseline and a fine-tuned version.
In addition to these CLMs, we implemented input encodings inspired by DeepJIT (diff with <ADD>/<DEL> tags) and JIT-Fine (addeddeleted block format). These were not used as standalone baselines, but rather as representation strategies integrated into our Stage 1 evaluation protocol.
4.3. Evaluation Metrics
We evaluate model performance using five standard classification metrics: Accuracy, Precision, Recall, F1-score(F1), and PR-AUC. Among these, F1 is adopted as the primary metric, as it balances precision and recall and is widely used for imbalanced datasets such as JIT-SDP (Hoang et al., 2019; Kamei et al., 2012; Hoang et al., 2020).
Accuracy measures the overall proportion of correctly classified modifications, but can be misleading under severe class imbalance.
Precision quantifies the fraction of predicted defective modifications that are truly defective, while Recall measures the fraction of actual defective modifications that are successfully identified.
F1 is the harmonic mean of precision and recall, and is widely regarded as the most informative single measure in imbalanced classification settings.
PR-AUC is the area under the Precision–Recall curve, which provides a threshold-independent view of the trade-off between precision and recall across all operating points. Note that for the decoder-based model (Qwen2.5), we primarily report the four discrete classification metrics, as calculating threshold-independent measures like PR-AUC is less straightforward due to its generative inference nature compared to encoder-based models with dedicated classification heads.
4.4. Implementation Details
To ensure the reproducibility and statistical significance of our results, we fine-tuned each CLM–encoding configuration ten times using different random seeds (12345–12354) and report the mean and standard deviation. All experiments were conducted using NVIDIA Tesla V100 (32GB) and RTX 3080 Ti (12GB) GPUs.
Encoder-based Training. CodeBERT, CodeT5+, and UniXcoder were fine-tuned for 6 epochs with a learning rate of , a batch size of 8, and the AdamW optimizer with decoupled weight decay (Loshchilov and Hutter, 2017). We employed a linear learning-rate scheduler with a 10% warm-up ratio and utilized gradient accumulation to stabilize training. To address class imbalance, we applied a cost-sensitive weighting scheme, assigning the positive class a weight of (King and Zeng, 2001). The best-performing model across the 6 epochs was selected based on validation F1 for final testing.
Decoder-based Training. For Qwen2.5-7B-Instruct, we utilized the Unsloth (Daniel Han and team, 2023) framework to perform memory-efficient Supervised Fine-Tuning (SFT) with 4-bit quantization. Given the large parameter scale, we applied Parameter-Efficient Fine-Tuning (PEFT) via LoRA () to the query, key, value, and projection layers. The model was trained for a single epoch with a learning rate of , an effective batch size of 16, and the 8-bit AdamW optimizer. To ensure optimal convergence within a single epoch, we performed intra-epoch validation at 10 equally spaced intervals and selected the checkpoint with the highest validation F1. Unlike the encoder models, class imbalance was addressed via oversampling the defective class to match the size of the clean class (Buda et al., 2018). The same prompt templates were utilized for both fine-tuning and zero-shot evaluations to ensure consistency, all of which are available in our replication package.
Statistical Analysis. For RQ2, we conducted a two-way repeated-measures ANOVA with within-subject factors model (6) and encoding (5), treating the 10 seeds as subjects. We report partial and Cohen’s as effect sizes (Lakens, 2013). Post-hoc comparisons used paired -tests with Holm step-down correction (Holm, 1979) applied per family: Encoding main-effect pairs on subject-level marginal means (collapsed over models), and all encoding pairs within each model (simple effects).
For RQ3, each perturbation was paired with its corresponding original encoding and evaluated across the same 10 seeds. We tested the directional hypothesis that perturbations degrade performance using one-sided paired -tests (), complemented by Wilcoxon signed-rank tests (Wilcoxon, 1945). We also report paired effect size as Cohen’s and non-parametric percentile 95% bootstrap confidence intervals (CIs) for the mean difference using 10,000 resamples (Tibshirani and Efron, 1993). -values were adjusted for multiple comparisons using Holm’s method within each set of tests (e.g., across the four perturbations). Following common conventions, we consider results statistically significant at . Effect sizes are interpreted using standard thresholds: partial (small), (medium), (large); Cohen’s (small), (medium), (large) (Cohen, 2013).
5. Experimental results
RQ1: How accurate are the labels in the ReDef JIT-SDP dataset?
We manually audited 100 randomly sampled defective functions. Following recent practices in vulnerability dataset curation (Chen et al., 2023; Ding et al., 2024), two authors independently inspected the commit evidence—reverted commit message, function_before, function_after, and revert commit message—and labeled each case as Correct, Ambiguous, or Wrong. After adjudication, 92 cases were classified as Correct, 5 as Ambiguous, and 3 as Wrong, yielding a positive-label precision of 0.92 as shown in Table 3. Compared to prior reports of 14–50% noise in SDP datasets (Herbold et al., 2022; Afric et al., 2023), our 8% mislabel rate is substantially lower.
| Correct | Ambiguous | Wrong | |
| Reviewer 1 | 94 | 5 | 1 |
| Reviewer 2 | 83 | 14 | 3 |
| Adjudicated | 92 | 5 | 3 |
Ambiguities initially accounted for 30% of cases, typically arising when revert messages cited only a hash without any rationale, forcing annotators to infer defect-relatedness. We revised the rubric to abstain unless defect links were explicit, which reduced ambiguous cases to 5% in the final adjudicated set. The three Wrong cases followed non-defect patterns (e.g., mistaken commit ordering, temporary mitigations, or structural refactoring).
Overall, these results suggest that our revert-anchored, history-screened pipeline yields comparatively reliable defect labels.
RQ2: Which input encodings yield the best CLM performance?
Table 4 summarizes the performance of all CLM–encoding combinations. Overall, Diff w/ tags and AddedDeleted consistently achieved the highest F1 across most models, while After+Markers and Before+After often performed worst, particularly in encoder-based models. Notably, we observed a recurring discrepancy where encodings with lower F1 often rank higher in accuracy. This discrepancy reflects the class imbalance in ReDef (3:1 clean-to-defective ratio), where accuracy is often inflated by a bias toward the majority clean class despite poor defect detection performance.
Beyond these task-specific adaptation results, Qwen2.5’s zero-shot performance exhibits a universal decline across all metrics. The After-only zero-shot configuration fails almost entirely (F1 0.1), underscoring that in zero-shot settings, providing explicit change information is crucial for the model to establish any predictive utility. To compare the fine-tuned models more rigorously beyond such raw patterns, we applied statistical tests as follows.
| Model | Encoding | F1 | PR-AUCa | Accuracy | Precision | Recall |
| CodeBERT | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| CodeT5+ | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| UniXcoder-512 | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| UniXcoder-1,024 | After-only | |||||
| After+Markers | ||||||
| Before+After | ||||||
| Diff w/ tags | ||||||
| AddedDeleted | ||||||
| Qwen2.5-512 | After-only | — | ||||
| After+Markers | — | |||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-1,024 | After-only | — | ||||
| After+Markers | — | |||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-512 | After-only | — | ||||
| (zero-shot) | After+Markers | — | ||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-1,024 | After-only | — | ||||
| (zero-shot) | After+Markers | — | ||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| aPR-AUC is omitted for Qwen2.5 due to its generative output nature. | ||||||
We conducted a two-way repeated-measures ANOVA with model (6 fine-tuned CLMs) and encoding (5 variants) as within-subject factors (Table 5). Results showed significant main effects of both model () and encoding (), both indicating large effect sizes. Post-hoc pairwise tests with Holm correction on the marginal means (collapsed over models) confirmed that Diff w/ tags and AddedDeleted significantly outperformed all other encodings (all ), while performing comparably to each other.
| Effect | df_num | df_den | ||||
| Model | 9.32 | 5 | 45 | 0.51 | 1.02 | |
| Encoding | 15.11 | 4 | 36 | 0.63 | 1.30 | |
| ModelEncoding | 1.91 | 20 | 180 | .014 | 0.17 | 0.46 |
Furthermore, a significant interaction effect was found (), suggesting that the impact of encoding strategies varies across different model architectures. Simple effect analysis revealed that for CodeBERT and CodeT5+, change-focused encodings (Diff w/ tags, AddedDeleted) consistently demonstrated a statistically significant performance advantage (). In contrast, for UniXcoder and Qwen2.5 models, the pairwise differences did not reach statistical significance. Nevertheless, a consistent numerical advantage of change-focused encodings was observed across all models, with Diff w/ tags or AddedDeleted consistently achieving the highest F1 (Table 4).
This consistent advantage can be explained by two primary factors: truncation risk and distracting context. First, as shown in Figure 3, only about 30–50% of functions fit within the 512-token limit. Consequently, whole-function encodings (After-only, After+Markers, Before+After) often truncate important edits, whereas compact diff-style formats preserve the critical change signal within the token budget. Second, supplying the entire function introduces excessive, mostly unchanged context that distracts the model from the actual modification. The weakness of Before+After illustrates both issues: two full functions must fit under the same budget, simultaneously increasing truncation risk and noise. Supporting this distracting context hypothesis, both UniXcoder and Qwen2.5 generally performed better under a 512-token limit than at 1,024 tokens. Therefore, the observed superiority of diff-style encodings is primarily a practical one; it reflects their efficiency in preserving critical signals within finite context windows rather than an inherent semantic advantage over whole-function views.
Interestingly, we observed an inverse correlation between model scale and performance. Despite having the largest parameter scale, Qwen2.5 exhibited the lowest overall performance. This trend extends to encoder-based models, where the model with the fewest parameters—CodeBERT—achieved the highest F1. Given that ReDef consists of difficult samples—modifications that professional developers initially committed and subsequently reverted—the signals for defects are likely extremely subtle. In such high-difficulty scenarios, larger models may struggle to discern meaningful features, while smaller models appear more adept at capturing simpler but consistent patterns.
While the absolute performance remains modest (F1 scores below 0.4), such results are common in classification tasks involving severe class imbalance and subtle signals, as characterized by the ReDef dataset (Guo et al., 2023; Ni et al., 2023). Nevertheless, as all models perform substantially above random baselines, they are evidently capturing discriminative patterns. This observation sets the stage for RQ3, where we investigate whether these learned patterns originate from a genuine understanding of code-change semantics or a reliance on shallow, invariant surface cues.
Finally, in our setting which primarily targets code understanding, CodeBERT again emerged as the strongest CLM, in line with recent studies showing its robustness on classification and comprehension tasks (Xiao et al., 2023; Dahou et al., 2025).
RQ3: Does CLM performance degrade under counterfactual perturbations of change representations?
To address RQ3, we conducted counterfactual perturbations using CodeBERT, the best-performing encoder-based model, and Qwen2.5-512, the strongest decoder-based variant. As shown in Table 6, across the four perturbations—Spurious change markers, Swapped snapshots, Reversed diff tags, and Swapped blocks—performance for both CodeBERT and Qwen2.5-512 remained effectively stable. Even when the logical direction or structural alignment of the modification was disrupted, F1 differences stayed within a negligible range (). This stability is counterintuitive: if these models were truly leveraging semantic signals from code changes, such distortions should have produced clear drops in performance metrics. Instead, the negligible differences suggest that the models are not semantically sensitive to modification content, but rather exploit surface-level regularities (e.g., token distributions, positional patterns) that persist under perturbation. In other words, the robustness observed here reflects insensitivity rather than genuine understanding.
The lack of response in the Swapped snapshots and Swapped blocks conditions ( after Holm correction for both models) provides compelling evidence of semantic blindness. Logically, reversing the states or blocks represents a functional opposite, yet both models were impervious to these changes. One might attribute this stability to the modest baseline performance; however, even a model with limited predictive power should exhibit a clear drop when its underlying causal logic is fundamentally reversed. Despite this consistent trend of insensitivity, we observed that Qwen2.5-512 exhibited slightly more sensitivity in certain conditions. Specifically, it showed a statistically significant decline under Reversed diff tags () and a marginal drop under Spurious change markers (). In contrast, CodeBERT remained almost entirely unaffected by these same perturbations, suggesting that encoder-based architectures may be even more reliant on shallow, invariant cues than their decoder-based counterparts. However, even for the larger decoder-based model, the absolute magnitude of these drops remained remarkably small, further confirming that its predictive utility still relies heavily on surface-level regularities that persist under perturbation. These results highlight a critical semantic gap: current CLMs achieve stability not through robust modeling of code-change logic, but by falling back on shallow features. This underscores the urgent need for architectures or training objectives specifically designed to capture the relational dynamics of code transformations.
Percent change is relative to the corresponding original encoding.
| Model | Condition | F1 | PR-AUC | Accuracy | Precision | Recall |
| CodeBERT | After+Markers | |||||
| Spurious CHG markers | ||||||
| +0.23% | ||||||
| CodeBERT | Before+After | |||||
| Swapped snapshots | ||||||
| +0.48% | ||||||
| CodeBERT | Diff w/ tags | |||||
| Reversed diff tags | ||||||
| CodeBERT | AddedDeleted | |||||
| Swapped added/deleted blocks | ||||||
| Qwen2.5-512 | After+Markers | — | ||||
| Spurious CHG markers | — | |||||
| Qwen2.5-512 | Before+After | — | ||||
| Swapped snapshots | — | |||||
| +0.30% | ||||||
| Qwen2.5-512 | Diff w/ tags | — | ||||
| Reversed diff tags | — | |||||
| Qwen2.5-512 | AddedDeleted | — | ||||
| Swapped blocks | — | |||||
| +0.17% |
6. Threats to Validity
6.1. Internal Validity
Subjectivity in Manual Verification. Determining whether a code modification is actually bug-inducing is an inherently difficult and subjective task. To mitigate this subjectivity, our manual audit process relied primarily on explicit commit-message evidence (e.g., statements such as “reverting due to a crash”) rather than a subjective interpretation of code logic. This is because such natural-language mentions provide the most direct and reliable justification from the developers who performed the revert. In cases where the commit messages were vague or lacked sufficient reasoning for the revert, we categorized them as incorrectly collected samples. Consequently, this conservative auditing approach revealed an error rate of only 8% (Table 3), thereby demonstrating the high precision of our curated dataset. Despite these rigorous criteria, human inspection is not infallible, and the potential for misjudgment during the verification of the 100 samples remains a threat to internal validity.
Input Encodings and Perturbation Strategies. We evaluated five representative encodings, including formats inspired by prior JIT-SDP studies. However, all encodings in our study are derived solely from function_before and function_after, which prevents us from leveraging richer semantic signals available in other sources such as commit messages. Moreover, even within these code-centric views, alternative ways of structuring the information—such as tree-structured or graph-based encodings—may provide richer signals of semantic change that are not fully captured by the linear, token-based representations used in this work. Additionally, we designed four perturbations as diagnostic probes specifically intended to assess semantic sensitivity. While our results demonstrated that models remain largely indifferent even under significant semantic disruptions, there may exist alternative perturbation strategies that could further expose the extent of this insensitivity or uncover different facets of modification reasoning. Exploring such alternatives and their alignment with diverse real-world scenarios remains an open avenue for future work.
Statistical Power. All statistical tests were based on 10 independent runs with different seeds. While this design controls stochastic variance, a larger number of repetitions could yield stronger statistical power and narrower confidence intervals. However, prior AI studies on code intelligence tasks rarely report extensive repetitions: many omit random seeds entirely or rely on a single run, and even the more rigorous works seldom exceed 10 repetitions due to the substantial computational cost of CLMs (Ye et al., 2023; Grishina et al., 2023). Running 30 or more trials per configuration is generally infeasible in practice. Our choice of 10 runs follows this common practice while still providing a reasonable trade-off between robustness and resource constraints.
6.2. External Validity
Dataset Coverage and Generalizability. We intentionally restricted the scope of ReDef to single-function modifications anchored by revert commits. Since reverts and single-function edits represent only a subset of all defective modifications, our dataset does not encompass the full spectrum of defect patterns, such as multi-function changes or bugs that are not easily identified for reversion. This focus may limit the generalizability of our findings to more complex, “tangled” commits. However, this conservative approach significantly mitigates the “blame-innocent” problem inherent in traditional SZZ-based labeling, ensuring high label precision. By prioritizing reliability over sheer volume, ReDef serves as a high-confidence benchmark that complements existing large-scale but noisy datasets, providing a necessary trade-off between coverage and predictive reliability.
Project Selection Bias. Our dataset was drawn from 22 large-scale open-source projects across diverse domains. However, the data volume is notably concentrated in a few massive repositories, particularly Linux and Chromium. This skew risks biasing our findings toward system-level software patterns, potentially limiting their generalizability to smaller applications or projects with different review and revert policies. Furthermore, industrial or proprietary projects could exhibit markedly different defect patterns, and including such settings in future work would provide stronger external validity.
Cross-Language Applicability. This study focused solely on C/C++ projects. Substantial syntactic and semantic differences across programming languages may limit the applicability of our findings. Extending evaluation to multi-language datasets is a key direction for future work. Such efforts should examine whether our observations hold consistently across different language families.
7. Conclusion
This work introduced ReDef, a high-confidence JIT-SDP dataset consisting of 3,164 defective and 10,268 clean function-level modifications. Manual audits confirmed a label precision of 92%, which is substantially higher than that of prior JIT-SDP corpora. Our curation pipeline—combining revert anchoring, longitudinal history screening, and GPT-assisted triage—effectively minimizes the label noise prevalent in existing datasets. By overcoming the limitations of conventional SZZ-based resources, ReDef provides a robust foundation for the research community to evaluate change-aware models. This benchmark not only enables reproducible comparisons but also facilitates the development of next-generation, edit-sensitive learning objectives.
Using ReDef, we evaluated four representative Code Language Models (CLMs)—CodeBERT, CodeT5+, UniXcoder, and Qwen2.5—across five input encoding strategies. Our results indicate that compact, change-focused encodings consistently achieve the best performance by emphasizing localized modifications and minimizing information loss due to truncation. Notably, we observed that larger context windows do not necessarily translate to better performance; both UniXcoder and Qwen2.5 performed worse under a 1,024-token budget than at 512 tokens, suggesting that excessive context can obscure critical change signals.
Counterfactual perturbations of change representations further revealed that CLMs’ performance remains effectively stable even under significant structural distortions. While encoder-based models were almost entirely unaffected by these perturbations, the decoder-based Qwen2.5 exhibited only marginal sensitivity in specific conditions. Even when we reversed the logical order of modifications or injected spurious change markers, F1 remained nearly identical. If these models had truly internalized the semantics of code edits, such fundamental distortions should have led to a precipitous decline in predictive performance.
This outcome exposes a critical blind spot in current evaluations of code-change models: what appears to be robustness is, in fact, a symptom of semantic blindness, where models fall back on surface-level cues rather than genuine understanding. Our findings caution against interpreting high benchmark scores as evidence of an advanced grasp of code evolution. Ultimately, this study underscores the need for new architectures and training objectives specifically designed to capture the relational and causal dynamics of code transformations.
Several threats to validity remain. While ReDef ensures high label precision, its focus on single-function reverts in large-scale C/C++ projects may limit generalizability to multi-function commits or other software domains. Furthermore, despite our manual audits, the inherent subjectivity of human inspection persists as a potential bias, and alternative encoding or perturbation strategies may exist that could uncover deeper semantic insights. Future research should explore the underlying causes of this semantic lag, potentially through edit-aware pre-training objectives or architectures specifically tailored for change reasoning. It is equally vital to investigate whether this reliance on superficial patterns persists in other semantics-sensitive tasks, such as vulnerability detection and automated code review. Finally, evaluating larger API-based models (e.g., GPT-4o) under the same protocol will reveal whether increased scale can effectively bridge the critical “semantic gap” exposed in this study.
We hope that ReDef and our evaluation protocol will serve as a foundation for the next generation of change-aware, semantics-sensitive defect prediction models.
Data Availability
An anonymous link containing the dataset, construction scripts, and experimental code is publicly available for reproducibility:https://figshare.com/s/4f202bc0921e26b41dc2.